OpenStudy (anonymous):

(6-y/y-2)-(8y-7/2-y)

OpenStudy (anonymous):

You want to have a common denominator before you can add. So write the lower term in the right expressions as -(y - 2) so you have an expression like: \[\ \frac{6 - y}{y - 2} \ + \ \frac{8y - 7}{y-2} \ = \ \frac{7y - 1}{y-2} \]

OpenStudy (anonymous):

The reason we have a plus sign is because we factored out a negative one in the denominator so that made it a plus sign.